# inst4] organization=IPTC, ETSI de Telecomunicación,addressline=Universidad Politécnica de Madrid, city=Madrid, postcode=28040, country=Spain

### v2

@@ F-1 Score: Provides a balanced measure combining Precision and Recall:
It is important to note that the comparison with single-GPU methods is only as fair as possible. We increase the training iterations of all single-GPU methods by a factor of 2 as much as possible on hardware with 24GB VRAM (a factor of 3 or more will cause most methods to GPU out-of-memory (OOM)), and all training parameters were adjusted accordingly. The full-scene results clearly indicate that our method achieves reconstructions with substantially greater detail and completeness than these baselines.
@@
Table 3 presents a quantitative evaluation comparing our proposed method with existing multi-GPU-based GS methods on the Matrix City dataset [25], and two additional custom-collected datasets (Scene1 and Scene2). Precision, recall, and F1-scores are reported at two different distance thresholds ($\tau=1.0m$ and $\tau=0.5m$).
Our method consistently achieves the best or second-best results across metrics, underscoring the effectiveness and superiority of our per-point partitioning strategy in overcoming scalability bottlenecks, enabling efficient and accurate reconstruction of complex urban scenes. CityGaussianV2 [34], another multi-GPU-based method, similarly requires initial coarse training on a single GPU followed by partitioning for multi-GPU fine-tuning. However, this training paradigm becomes infeasible for datasets comprising nearly 10,000 images, such as Scene1 and Scene2, where the initial sparse point cloud generated by COLMAP exceeds 30 million points, surpassing the memory limitations of a single RTX 4090 GPU with 24GB VRAM. To address this challenge, we applied our proposed per-point partitioning strategy, dividing these two extensive scenes into smaller, manageable 2×2 sub-regions to apply the standard CityGaussianV2 training strategy. Consequently, as shown in the Table 3, this adaptation significantly improves the performance of CityGaussianV2 on the larger scenes.
@@ 5.6.1 Impact of Removing Individual Constraints
When both multi-view depth and normal consistency constraints are removed, it exhibits a noticeable performance decline, reinforcing the importance of geometric constraints in high-quality reconstruction. Interestingly, the F1 score for this condition is slightly higher than that of removing normal consistency alone (Condition C). This suggests that when only normal consistency is absent, the retained depth consistency may introduce conflicting geometric that worsen reconstruction errors, whereas removing both constraints forces the model to rely solely on single image-based information, leading to a different but somewhat more stable degradation. Figure 12 (a) and (b) further highlights the dominant role of multi-view depth and normal consistency in preventing surface misalignment and ensuring smooth reconstructions.
@@
When the densification process is removed, Precision and Recall drop slightly to 0.678 and 0.685, respectively, with an F1 Score of 0.682. However, the qualitative comparison in Figure 12 and 13 suggests that densification primarily refines the reconstruction rather than fundamentally altering its structure. The results indicate that while densification improves point distribution and mesh completeness, the geometric consistency constraints (depth and normal consistency) have a more significant impact on the overall reconstruction quality. When remove window mask, ULSR-GS runs out of memory (OOM), preventing evaluation. This suggests that removing the window mask leads to an excessive increase in Gaussian points, which significantly inflates the computational burden. The failure to complete the reconstruction in this setting highlights the importance of an adaptive sparsification mechanism, which prevents unnecessary computational overhead while preserving reconstruction quality.
@@ 5.6.3 Impact of Partition Strategies
Notably, Figure 15 visually illustrates the effectiveness of our partitioning strategy using a more challenging close-range oblique photogrammetry dataset featuring intricate and specific UAV flight paths [35]. In this dataset, our approach successfully identifies and selects images highly relevant to the reconstruction region, significantly enhancing local detail preservation and accuracy. Conversely, the partitioning strategy of Vast Gaussian, which relies strictly on the projection of camera positions onto a 2D plane, selects images from a broader area without adequate precision. Consequently, critical images positioned outside of a predefined boundary, yet crucial for reconstructing complex structures, are erroneously excluded. Simultaneously, the inclusion of unrelated, distant images introduces noise, ultimately deteriorating the reconstruction quality. Additionally, on datasets collected using multi-directional Penta Cam setups in Figure 14, where camera paths and orientations are highly irregular, our point-to-photo partitioning approach consistently outperforms the camera projection-based method.
@@
Table 7 quantitatively compares different partitioning methods, specifically our point-based partitioning approach and the image-based method employed by Vast Gaussian [29]. The results indicate that the Vast Gaussian partitioning strategy is more suitable for rendering tasks. This is because our method prioritizes the reconstruction of the scene’s primary structure by excluding the distant sparse SfM points during sub-region partitioning. Figure 16 provides a qualitative comparison to further illustrate this difference. When training ULSR-GS using different partitioning strategies, our partition method shows a trade-off: while it loses performance in reconstructing the background, it preserves significantly more detail in the reconstruction of the primary scene components.
@@ 6.1 Real World Application: Integration with USV Bathymetry model
We now discuss the ability of integrating the bathymetry digital elevation model (BDEM) obtained from single-beam echo sounder (SBES) with the oblique photogrammetric models produced by ULSR-GS to generate a real urban surface model. As shown in Figure 17 (a), the experimental site for this study is located at Moon Bay (Scene 2), situated along the southern shore of Dushu Lake in Suzhou, China. Moon Bay is a scenic and ecologically rich area, but it is also surrounded by a dense cluster of urban buildings, which presents significant challenges for UAV-based photogrammetric reconstruction and the subsequent data fusion with USV-derived bathymetric models.
@@
6.2 Strength and Limitations
Our proposed ULSR-GS framework demonstrates notable advantages and specific application scenarios that clearly distinguish it from existing single-GPU [15, 16, 66, 68, 5], multi-GPU [34, 30], and MVS-based [44] methods. Compared to single-GPU Gaussian Splatting (GS) approaches, ULSR-GS exhibits exceptional scalability and efficiency, making it particularly suitable for large-scale urban reconstruction tasks. The inherent memory constraints of single-GPU methods severely limit their capacity for handling extensive datasets; our multi-GPU strategy circumvents this limitation by employing an optimized point-to-photo partitioning technique driven by SfM. Consequently, ULSR-GS effectively reconstructs detailed urban scenes at scales previously unattainable with single-GPU methods [15, 16, 66, 68, 5].
Furthermore, compared with other multi-GPU GS approaches [34, 30] that typically rely on image-centric partitioning, ULSR-GS provides greater adaptability and accuracy, especially beneficial in complex scenarios such as close-range and oblique photogrammetry. Our innovative per-point partitioning ensures that each sub-region is reconstructed using its most relevant image set, thus addressing critical issues of uneven coverage and irregular camera paths encountered by traditional methods. These improvements are validated by quantitative metrics and qualitative evaluations across diverse urban scenes.
Moreover, relative to MVS-based methods [44], ULSR-GS is particularly effective in reconstructing thin and complex architectural structures, which often present significant challenges for correspondence-based MVS techniques due to matching ambiguities or limited local texture.
Despite these strengths, ULSR-GS also exhibits several limitations: (1) As shown in Figure 10, the framework is less effective in vegetation-rich environments, where dense correspondence-based MVS techniques better capture the intricate and semi-transparent structures inherent in foliage. (2) The generated meshes tend to be overly smooth, potentially lacking essential high-frequency geometric details needed for accurately representing intricate textures and detailed surfaces. (3) The current GS-based training and rendering resolution limit (approximately 1.6K) is insufficient for applications demanding high-detail fidelity, such as high-resolution (2cm per pixel) oblique photogrammetry. (4) Additionally, the existing depth-projection-based densification strategy primarily enhances regions already adequately reconstructed, inadequately addressing areas that are initially sparse or poorly captured.
@@ 7 Conclusion
We present ULSR-GS, a framework dedicated to high- fidelity surface extraction in ultra-large-scale scene. Specifically, our partitioning approach combines with a multi-view selection strategy. Additionally, ULSR-GS employs a multi-view geometric consistency densification to enhance surface details. Experimental results demonstrate that ULSR-GS outperforms other SOTA GS-based works on large-scale benchmark datasets. Our approach addresses key limitations of image-based partitioning methods [29, 33, 20] by ensuring finer granularity in mesh boundaries and prioritizing the reconstruction of key scene regions. This strategy enables us to extract meshes individually rather than merge sub-regions and extract the full scene. Unlike existing MVS-based approaches, our method enables the simultaneous generation of a high-quality radiance field for rendering and a detailed mesh representation for reconstruction, all within a single training pipeline. Also, comparison with traditional MVS-based methods demonstrates ULSR-GS drastically reduces processing time while maintaining fine structural details. These advantages make ULSR-GS a promising choice for oblique photogrammetry, offering an efficient and scalable solution for real-world applications.

### v3

@@ F-1 Score: Provides a balanced measure combining Precision and Recall:
The comparison with single-GPU methods is conducted on the GauU-Scene [54] dataset. It is still important to note that the comparison with these methods is only as fair as possible. We increase the training iterations of all single-GPU methods by a factor of 2 as much on hardware with 24GB VRAM (a factor of 3 or more will cause most methods to GPU out-of-memory (OOM)), and all training parameters were adjusted accordingly. The full-scene results clearly indicate that our method achieves reconstructions with substantially greater detail and completeness than these baselines.
@@
Table 3 presents a quantitative evaluation comparing our proposed method with existing multi-GPU-based GS methods on the Matrix City dataset [25], and two additional custom-collected datasets contain nearly 10,000 images (Scene 1 and Scene 2). Precision, recall, and F1-scores are reported at two different distance thresholds ($\tau=1.0m$ and $\tau=0.5m$).
Our method consistently achieves the best or second-best results across metrics, underscoring the effectiveness and superiority of our per-point partitioning strategy in overcoming scalability bottlenecks, enabling efficient and accurate reconstruction of complex urban scenes. CityGaussianV2 [34] requires initial coarse training on a single GPU followed by partitioning for multi-GPU fine-tuning. However, this training paradigm becomes infeasible for datasets comprising nearly 10,000 images, such as Scene1 and Scene2, where the initial sparse point cloud generated by COLMAP exceeds 30 million points, surpassing the memory limitations of a single RTX 4090 GPU with 24GB VRAM. To address this challenge, we applied our proposed per-point partitioning strategy, dividing these two extensive scenes into smaller, manageable 2×2 sub-regions to apply the standard CityGaussianV2 training strategy. Consequently, as shown in Table 3, this adaptation significantly improves the performance of CityGaussianV2 on the larger scenes.
@@ 5.6.1 Impact of Removing Individual Constraints
When both multi-view depth and normal consistency constraints are removed, it exhibits a noticeable performance decline, reinforcing the importance of geometric constraints in high-quality reconstruction. Interestingly, the F1 score for this condition is slightly higher than that of removing normal consistency alone (Condition C). This suggests that when only normal consistency is absent, the retained depth consistency may introduce conflicting geometry that worsens reconstruction errors, whereas removing both constraints forces the model to rely solely on single image-based information, leading to a different but somewhat more stable degradation. Figure 12 (a) and (b) further highlight the dominant role of multi-view depth and normal consistency in preventing surface misalignment and ensuring smooth reconstructions.
@@
When the densification process is removed, Precision and Recall drop slightly to 0.678 and 0.685, respectively, with an F1 Score of 0.682. However, the qualitative comparison in Figure 12 and 13 suggests that densification primarily refines the reconstruction rather than fundamentally altering its structure. The results indicate that while densification improves point distribution and mesh completeness, the geometric consistency constraints (depth and normal consistency) have a more significant impact on the overall reconstruction quality. When removing the window mask, ULSR-GS runs out of memory (OOM), preventing evaluation. This suggests that removing the window mask leads to an excessive increase in Gaussian points, which significantly inflates the computational burden. The failure to complete the reconstruction in this setting highlights the importance of an adaptive sparsification mechanism, which prevents unnecessary computational overhead while preserving reconstruction quality.
@@ 5.6.3 Impact of Partition Strategies
Notably, Figure 15 visually illustrates the effectiveness of our partitioning strategy using a more challenging close-range oblique photogrammetry dataset featuring intricate and specific UAV flight paths [35]. In this dataset, our approach successfully identifies and selects images highly relevant to the reconstruction region, significantly enhancing local detail preservation and accuracy. Conversely, the partitioning strategy of Vast Gaussian, which relies strictly on the projection of camera positions onto a 2D plane, selects images from a broader area without adequate precision. Consequently, critical images positioned outside of a predefined boundary, yet crucial for reconstructing complex structures, are erroneously excluded. Simultaneously, the inclusion of unrelated, distant images introduces noise, ultimately deteriorating the reconstruction quality. Additionally, on datasets collected using multi-directional Penta-Cam setups in Figure 14, where camera paths and orientations are highly regular, our point-to-photo partitioning approach consistently outperforms the camera projection-based method.
@@
Table 7 quantitatively compares different partitioning methods, specifically our point-based partitioning approach and the image-based method employed by Vast Gaussian [29]. The results indicate that the Vast Gaussian’s [29]] partitioning strategy is more suitable for rendering tasks. This is because our method prioritizes the reconstruction of the scene’s primary structure by excluding the distant sparse SfM points during sub-region partitioning. Figure 16 provides a qualitative comparison to further illustrate this difference. When training ULSR-GS using different partitioning strategies, our partition method shows a trade-off: while it loses performance in reconstructing the background, it preserves significantly more detail in the reconstruction of the primary scene components.
@@ 6.1 Real World Application: Integration with USV Bathymetry model
We now discuss the ability to integrateg the bathymetry digital elevation model (BDEM) obtained from a single-beam echo sounder (SBES) with the oblique photogrammetric models produced by ULSR-GS to generate a real urban surface model. As shown in Figure 17 (a), the experimental site for this study is located at Moon Bay (Scene 2), situated along the southern shore of Dushu Lake in Suzhou, China. Moon Bay is a scenic and ecologically rich area, but it is also surrounded by a dense cluster of urban buildings, which presents significant challenges for UAV-based photogrammetric reconstruction and the subsequent data fusion with USV-derived bathymetric models.
@@
6.2 Strengthsh and Limitations
Our proposed ULSR-GS framework demonstrates notable advantages and specific application scenarios that clearly distinguish it from existing single-GPU [15, 16, 66, 68, 5], multi-GPU [34, 30], and MVS-based [44] methods. Compared to single-GPU GS approaches, ULSR-GS exhibits exceptional scalability and efficiency, making it particularly suitable for large-scale urban reconstruction tasks. The inherent memory constraints of single-GPU methods severely limit their capacity for handling extensive datasets; our multi-GPU strategy circumvents this limitation by employing an optimized point-to-photo partitioning technique driven by SfM. Consequently, ULSR-GS effectively reconstructs detailed urban scenes at scales previously unattainable with single-GPU methods [15, 16, 66, 68, 5].
Furthermore, compared with other multi-GPU GS approaches [34, 30] that typically rely on image-centric partitioning, ULSR-GS provides greater adaptability and accuracy, especially beneficial in complex scenarios such as close-range and oblique photogrammetry. Our per-point partitioning ensures that each sub-region is reconstructed using its most relevant image set, thus addressing critical issues of uneven coverage and irregular camera paths encountered by traditional methods. These improvements are validated by quantitative metrics and qualitative evaluations across diverse urban scenes. Moreover, relative to MVS-based methods [44], ULSR-GS is particularly effective in reconstructing thin and complex architectural structures, which often present significant challenges for correspondence-based MVS techniques due to matching ambiguities or limited local texture.
Despite these strengths, ULSR-GS also exhibits several limitations: (1) As shown in Figure 10, the framework is less effective in vegetation-rich environments, where dense correspondence-based MVS techniques better capture the intricate and semi-transparent structures inherent in foliage; (2) the current GS-based training and rendering resolution limit (approximately 1.6K) is insufficient for applications demanding high-detail fidelity, such as high-resolution (2cm per pixel) oblique photogrammetry; and (3) the existing depth-projection-based densification strategy primarily enhances regions already adequately reconstructed, inadequately addressing areas that are initially sparse or poorly captured.
@@ 7 Conclusion
We present ULSR-GS, a framework dedicated to high- fidelity surface extraction in ultra-large-scale scenes. Specifically, our partitioning approach combines with a multi-view selection strategy. Additionally, ULSR-GS employs a multi-view geometric consistency densification to enhance surface details. Experimental results demonstrate that ULSR-GS outperforms other SOTA GS-based works on large-scale benchmark datasets. Our approach addresses key limitations of image-based partitioning methods [29, 33, 20] by ensuring finer granularity in mesh boundaries and prioritizing the reconstruction of key scene regions. This strategy enables us to extract meshes individually rather than merge sub-regions and extract the full scene. Unlike existing MVS-based approaches, our method enables the simultaneous generation of a high-quality radiance field for rendering and a detailed mesh representation for reconstruction, all within a single training pipeline. Also, comparison with traditional MVS-based methods demonstrates ULSR-GS drastically reduces processing time while maintaining fine structural details. These advantages make ULSR-GS a promising choice for oblique photogrammetry, offering an efficient and scalable solution for real-world applications.